OpenStudy (anonymous):

If $ 2500 is invested in an account that pays interest compounded continuously, how long will it take to grow to $ 5000 at 3%? Round to the nearest tenth

OpenStudy (anonymous):

this is a log questionExponential & Logarithmic Functions

OpenStudy (anonymous):

The equation would be 2500*1.03^x=5000

OpenStudy (anonymous):

1.03^x=5000/2500=2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but I taught log where simplying the equation or i'm wrong

OpenStudy (anonymous):

You have to solve x... (I didn't finish :p)